Magical Mathematics[Interesting Approach]

Here is a question of speed ,time and distance...

The speed of a motor boat itself is 20 km/h and the rate of flow of the river is 4 km/h. Moving with the stream the boat went 120 km. What distance will the boat cover during the same time going against the stream?

rekha arora
17 Years agoGrade

1 Answer

Jitender Pal
17 Years ago

Relative speed towards the stream of a boat

(20km/h+4km/h) X ( time)=120km

time= 120/24 h

time= 5 hrs

now relative speed against stream

(20-4)km/h X time = distance

16 km/h X 5 = distance

80 km